Please help w/ the following:
M(6,-5) is the midpoint of RS. If S has coordinates (14,0), find the coordinates of R.
Thank you.

Answer by venugopalramana(3286)

Please help w/ the following:
M(6,-5) is the midpoint of RS. If S has coordinates (14,0), find the coordinates of R.
Thank you.
FORMULA FOR COORDINATES OF MID POINT (X,Y) BETWEEN (X1,Y1) AND (X2,Y2) IS
X=(X1+X2)/2.....Y=(Y1+Y2)/2.SO
6=(14+X2)/2..........-5=(0+Y2)/2
12=14+X2............-10=Y2
X2=-2............Y2=-10
R IS (-2,-10)